Given that x and y are real numbers satisfying the equations 3x−2y=21 and 7x+3y=95, what is the value of x+y?
Answer by sofiyac(983)

STEP 1:
Multiply each equation by a number that will create
opposite coefficients for x.
In this case we will multiply first equation
by -7 and the second equation by 3;
After multiplying we have the following system:
-21x+14y=-147
21x+9y-285


STEP 2: add the two equations together to eliminate x from the system.
23y=138

STEP 3: find y
y=6
STEP 4: substitute the value for y into the original equation to solve for x.
3x-2(6)=21
3x=33
x=11

add x+y and you get 11+6=17
Sorry about that, must have made a mistake somewhere!!!! :)